OpenStudy (anonymous):

integral from 0 to x e^t^2 use Taylor or Maclaurin series to solve

OpenStudy (anonymous):

Well, the Maclaurin series of e^x is \[\sum_{n=0}^{\infty}x^n/n!\] You can replace x with t^2.

OpenStudy (anonymous):

Ohp there was a multiple choice part...hold on let me pull that up

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

a.) summation x^2n/n factorial b.) e^x^2-1 c.) x^(2n+1)/(2n+1)nfactorial d.) x^2n+1/(n+1) factorial e.) x^(2n+1)/(n+1)(2n)factorial

OpenStudy (anonymous):

c d and e have summation also

OpenStudy (anonymous):

I got c or: \[\int\limits_{0}^{x} \sum_{n=0}^{\infty} t^{2n}/n! dt=\sum_{n=0}^{\infty}t^{2n+1}/(2n+1)n! |_0^x\]

OpenStudy (anonymous):

Or \[\sum_{n=0}^{\infty}x^{2n+1}/(2n+1)n! \] Your answer choice c

OpenStudy (anonymous):

Oooo I see

OpenStudy (anonymous):

Thank you sir!

OpenStudy (anonymous):

I have one more question but I should go a new one so I can give you another medal lol

OpenStudy (anonymous):

Okay :) Thank you

OpenStudy (anonymous):

no thank you :) alright let me post it real quick...